If the mass of a balloon is 2.0 g, its coefficient of static friction with the wall is 0.79, and the average distance between the opposite charges is 0.75 mm, what minimum amount of charge must be placed on the balloon in order for it to stick to the wall?
Answer #1

in order to stickthe balloon with wall

fricitional force = gravitational force on balloon

u F = mg

u k q^2 / r^2 = mg

0.79 * 9*10^9* q^2 / (0.75*10^-3)^2 = 2*10^-3* 9.8

q = 1.245*10^-9 C

=========

Comment in case any doubt, will reply for sure.. Goodluck